Correct combination of numbers and unit/substances from the drop down menus to setup and solve the following problem


how many grams are in 5.32 x 10(24) atoms of neon


1
Expert's answer
2020-11-09T13:50:25-0500

A mole of neon atoms (with the formula Ne) contains 6.02×1023 Ne atoms.

Proportion:

1 mol - 6.02×1023 Ne atoms

x mol - 5.32×1024 Ne atoms

"x = \frac{5.32 \times 10^{24}}{6.02 \times 10^{23}}"

x = 8.83 mol

M(Ne) = 20.18 g/ mol

m = n×M

m(Ne) = 8.83 × 20.18 = 178.19 g

Answer: 178.19 g